OpenStudy (anonymous):

Last Question of the day The function f is continuous on the interval [4, 15], with some of its values given in the table above. Estimate the average value of the function with a Trapezoidal Approximation, using the 4 intervals between those given points. x 4 9 11 14 15 f(x) –6 –11 –18 –21 –25

OpenStudy (anonymous):

okay so i'v done this a couple of times but i keep getting the wrong answer.

OpenStudy (solomonzelman):

you have a bunch of trapezoids, that is all. Area of each trapezoid is: \(\large\displaystyle A=(\Delta x)(h_1+h_2)/2\) where \(\Delta x\) is the width and \(\large\displaystyle (h_1+h_2)/2\) is the average height \(\large\displaystyle Trapezoid_1=(9-4)(-6+-11)/2\) \(\large\displaystyle Trapezoid_2=(11-9)(-18+-11)/2\) \(\large\displaystyle Trapezoid_3=(14-11)(-21+-18)/2\) \(\large\displaystyle Trapezoid_4=(15-14)(-25+-21)/2\)

OpenStudy (solomonzelman):

then add all trapezoids.

OpenStudy (solomonzelman):

Dont wonder the area is negative, that is because the entire thing is under the x-axis.

OpenStudy (anonymous):

yep im getting the same answer i was getting before , still wrong

OpenStudy (anonymous):

these are my answer choices –12.727 –11.546 –16.273 –13.909

OpenStudy (anonymous):

@SolomonZelman you still there ?

OpenStudy (solomonzelman):

yes, i am here. I am just glitching outside the US a little.

OpenStudy (solomonzelman):

what did you get when you added the trapezoids?

OpenStudy (anonymous):

153, i did the same thing but by taking the left point and right point sum and averaging them

OpenStudy (solomonzelman):

Yes, the area of the trapezoid is given by the product of its average height times the width, and the area of your full shape is the sum of all trapezoids... i dont understand why these options are here.

OpenStudy (solomonzelman):

i get -153 doing a completely valid method.... maybe your table has typos?

OpenStudy (anonymous):

perhaps we're finding the wrong thing . it says find the average value might mean the average slope not average area under the curve

OpenStudy (anonymous):

@ganeshie8 any idea?

OpenStudy (freckles):

\[\frac{1}{b-a} \int\limits _a^b f(x) dx\]

OpenStudy (joannablackwelder):

The average function value formula is: